# Photon distillation schemes with reduced resource costs based on multiphoton Fourier interference

### v2

@@ Abstract
Improving the indistinguishability of single photons is a crucial prerequisite for achieving large-scale photonic quantum computation. Photon distillation uses quantum interference to enhance the quality of single photons, sacrificing multiple photons to generate one photon with enhanced indistinguishability. By studying multiphoton interference in Fourier matrices, we find photon distillation schemes that require fewer photons to achieve the same improvement in indistinguishability, compared to the state of the art. These results may find application as a component in large-scale photonic quantum computers.
@@ I Introduction
Achieving quantum computation beyond the NISQ era requires large-scale, error-corrected devices [1]. Progress towards fault-tolerant quantum computing is achieved both by reducing physical noise and by improving quantum-error correction schemes to relax threshold requirements. However, physical noise mechanisms are hardware-dependent, thereby requiring mitigation strategies specific to a particular computational platform.
@@
Since the introduction of the Knill-Laflamme-Milburn scheme, linear optics stands out as a promising hardware platform for fault-tolerant photonic quantum computing [2]. In linear optics, allowable operations are restricted to measurement, linear optical circuits, i.e. optical devices which act linearly on the electric field amplitude, and classical reconfiguration of such circuits based on earlier measurement outcomes. Modern efforts center on fusion-based quantum computing (FBQC) [3, 4, 5], a computational architecture where multiple fixed-size entangled states - resource states - are entangled by fusion measurements. Recent experimental progress, for example in experimentally generating noisy three-qubit entangled states [6, 7, 8], shows the potential of this protocol.
@@
The key resource for all linear-optical quantum computing schemes is photon indistinguishability [9, 10, 11], i.e. the degree to which the photons in the quantum computer are identical in the degrees of freedom of those photons which are not being used to encode information. The reason for the importance of indistinguishability as a resource is that while linearity forbids an explicit photon-photon interaction, it allows for multi-particle quantum interference, which is sufficient for universal quantum computation when combined with measurements and fast feedback. The archetypal example of multi-particle interference is the Hong-Ou-Mandel effect [12]. Since photon indistinguishability governs the degree to which multi-particle interference occurs, it is the crucial resource for photonic quantum computing in its modern implementations [13, 14, 15, 16].
@@
Photon distillation works by injecting multiple photons into a linear optical circuit and performing a measurement on some faction of the outputs of that circuit. Conditional on observing a pre-defined measurement outcome (the herald) in all but one of the outputs of the circuit, the remaining output is projected onto a state with improved photon indistinguishability. To improve the degree of indistinguishability of photons, photonic distillation schemes [17, 18] have been proposed and experimentally demonstrated [19]. Photon distillation (see Fig. 1) is a nondeterministic method for heralding indistinguishable photons. These schemes trade numbers for quality, exchanging multiple weakly indistinguishable photons for one photon with reduced distinguishability error, analogous to magic state distillation [20]. However, in contrast to conventional filtering, e.g. spectrally, photon distillation does not require prior knowledge about the target wave function of the indistinguishable photons. Moreover, photon distillation filters in all internal degrees of freedom simultaneously.
The resource cost $C$ of such schemes is given by the number of photons of a given degree of indistinguishability that is on average required to obtain one photon of a given improved indistinguishability, or equivalently, the degree of improvement that can on average be achieved with a fixed number of photons. This resource cost, in turn, is determined by two factors: the error reduction $R$, i.e. the degree to which a single successful application of the scheme reduces the indistinguishability error on each photon, and the heralding probability $p$, i.e. the probability with which the scheme succeeds in the first place. In the limit of low errors, the error reduction reduces to a multiplicative factor, i.e. $\epsilon^{\prime}=\epsilon/R$, where $\epsilon$ and $\epsilon^{\prime}$ are the degree of error before and after the protocol, respectively. The probabilistic nature of the scheme can be mitigated by multiplexing [21], and for schemes where the available error reduction is limited, further error reduction is available by concatenation, i.e. using the input from one distillation step as input for the next. The current state-of-the-art photon distillation protocol is presented in Ref. [18], with a degree of error reduction of $R=3$ and a heralding success probability $p=\frac{1}{3}$. This scheme requires concatenation and therefore has a resource cost quadratic in the overall error reduction required.
In this work, we demonstrate distillation methods for single photon state purification that have a lower resource cost than the state-of-the-art. We recognize that the scheme proposed in [18], originally found by numerical optimization, in fact implements a Fourier transform on the optical modes. Following on this observation, we both numerically and theoretically study Fourier-based distillation schemes with arbitrary numbers of photons, using simulation techniques originally designed for boson sampling [10, 22, 23, 24, 25]. We find that larger Fourier matrices give rise to more resource-efficient distillation schemes, compared to the scheme of [18]. This is due to the fact that for 5-mode Fourier transformations and larger, multiple detection outcomes herald a successful distillation step. Up to a conjecture in the theory of permanents of Fourier matrices, we find strong evidence that the probability of heralding success tends to a constant as a function of system size. This, combined with improved error reduction at larger system sizes, removes the need for concatenation, changing our resource cost from quadratic in the desired error reduction to linear. Besides studying the scaling in the limit of large numbers of photons, we discuss how our results can be implemented in systems of finite size. Our work contributes to the design of fault-tolerant photonic quantum computing architectures by cutting down resource costs for fault-tolerant photonic quantum computing.
Note added. - Related work was simultaneously published [26].
@@
Figure 1: Artist’s impression of photon distillation. Partially distinguishable photons emitted by imperfect sources are interfered by a photonic distillation circuit. A given set of partial measurements heralds a quantum state with increased partial distinguishability. After a successful distillation attempt, this purified photon is ready to be processed for fusion based quantum computing. The distillation step reduces physical noise levels, which is important to achieve quantum error correction thresholds.
@@ III Fourier transform-based distillation
In the previous section, we established the framework for imperfect photon interference. To establish intuition for the problem of purification in Fourier matrices, we first extensively study the $N=5$ photon Fourier transform with a single photon at each input mode, in both error reduction and heralding probability. A detailed description of our numerical simulation methods is presented in App. B.
Fig. 2 shows the error reduction as a function of input error, for different partial measurement outcomes on modes 1 through 4. Surprisingly, there is more than one valid herald measurement for successful photon distillation. We observe three distinct families of successful heralding patterns.
@@ IV Theoretical analysis
$p=\sum_{j=0}^{N}(-1)^{j}(j+1)\prod_{i=1}^{j}(1-\frac{i}{N}).$ (14)
@@
$p=\frac{\frac{1}{N}\binom{2N-3}{N-1}}{\frac{1}{N}\binom{2N-1}{N}}=\frac{1}{4-\frac{2}{N}}=\frac{1}{4}+\frac{1}{8N}+\mathcal{O}((\frac{1}{N})^{2}).$ (15)
@@ V Discussion
VI Conclusion
@@
In conclusion, we have used a combination of numerical and theoretical tools to shown that generalized Fourier transform-based distillation schemes can be used to purify photons from imperfect sources. For this purpose, we explicitly simulated the error reduction for the $N=5$ photon Fourier transform where we showed that there are different paths towards perfect photon indistinguishability. Within the limit of low partial distinguishability error, we have shown modulo a conjecture that the error reduction scales linearly with the number of photons employed in the distillation scheme $N$, while we have given numerical and theoretical evidence that the success probability tends to a constant, giving rise to linear scaling in the number of photons required to achieve a given error reduction.
@@
Applicability of our work strongly depends on the availability of high-fidelity programmable photonic processors and fast optical switches. However, we note that processor size and fidelity of linear optical circuits are constantly improving [31, 44, 45, 46, 47, 48]. Our generalized schemes contribute to the development of adaptive purification circuits, where concatenation strategies are dependent on the actual herald measurements of each distillation step. While current advances in single photon source engineering continue to improve photon indistinguishability [49, 50, 51, 52], we anticipate that photon distillation will aid in the task of creating the millions of almost fully indistinguishable photons necessary for photonic quantum computing.
@@ Appendix C Multi-photon interference coefficients
For comparison, we inspected the $c_{j}$ distribution for Haar-random unitary matrices. We generate $10^{4}$ Haar-random matrices and compute the ratio $\sum_{j=0}^{N}c_{j}j$ over $\sum_{j=0}^{N}c_{j}$ for all mode-assignment lists with exactly one in mode $N$. The histogram of this ratio, for $N=3$ and $N=5$ is plotted in Fig. 9. We observe that the herald measurements for the Fourier transforms (dashed and dotted lines) have extreme ratios that satisfy Eq. 10.

### v3

@@ Abstract
We present a scalable scheme to achieve photon distillation, i.e. the preparation of a single photon with reduced indistinguishability error out of multiple imperfect photons, based on multiphoton interference in Fourier matrices. Our scheme achieves arbitrary error reduction in a single step, removing the need to concatenate multiple rounds of the protocol. Our scheme has resource costs linear in the error reduction, and a constant probability of success for arbitrary error reductions. We find that the scheme both raises the threshold of error correcting codes in photonic quantum computers, as well as reduces the resources required to construct a photonic quantum computer, making it highly promising for implementation in such devices.
@@ I Introduction
Since the introduction of the Knill-Laflamme-Milburn scheme, linear optics stands out as a promising hardware platform for universal, fault-tolerant photonic quantum computing [1], due to the long coherence time of photons and the high-precision control available over optical quantum states. In linear optics, the allowable operations are restricted to measurement, linear optical circuits, i.e. optical devices which act linearly on the electric field amplitude, and classical reconfiguration of such circuits based on earlier measurement outcomes. Modern efforts center on fusion-based quantum computing (FBQC) [2, 3, 4], a computational architecture where multiple fixed-size entangled states - resource states - are entangled by fusion measurements. Recent experimental progress, for example in experimentally generating noisy three-qubit entangled states [5, 6, 7], shows the potential of this protocol.
@@
The key resource for all linear-optical quantum computing schemes is photon indistinguishability [8, 9, 10], i.e. the degree to which the photons in the quantum computer are identical in the degrees of freedom of those photons which are not being used to encode information. Photonic indistinguishability is essential because while linearity forbids an explicit photon-photon interaction, it allows for multi-particle quantum interference, which is sufficient for universal quantum computation when combined with measurements and fast feedback. The archetypal example of such interference is the Hong-Ou-Mandel effect [11]. Since photon indistinguishability governs the degree to which multi-particle interference occurs, it is the crucial resource for photonic quantum computing in its modern implementations [12, 13, 14, 15], and imperfect indistinguishability gives rise to logical errors in the computation.
@@
To improve the degree of indistinguishability of photons, photonic distillation schemes [16, 17] have been proposed and experimentally demonstrated [18]. Photon distillation (see Fig. 1) is a nondeterministic method for heralding photons with a reduced indistinguishability error. These schemes trade numbers for quality, using interference and measurement to exchange multiple weakly indistinguishable photons for one photon with reduced distinguishability error, analogous to magic state distillation [19]. In contrast to optical filtering, photon distillation does not require prior knowledge about the target wave function of the indistinguishable photons, and filters in all internal degrees of freedom simultaneously. However, existing photon distillation protocols [16, 17] suffer from substantial resource costs due to the fact that each round of the protocol only moderately reduces the indistinguishability error, requiring concatenation to achieve higher error reduction. Since each round fails with a constant probability, this requires multiplexing [20] and intermediate optical feed-forward, which results in resource requirements which render the protocol infeasible in practice.
In this work, we present a photon distillation scheme with significantly reduced resource costs, rendering the protocol feasible for implementation in photonic quantum computers. Our scheme is based on multiphoton interference in Fourier matrices. In contrast to the state of the art [17], our scheme can achieve arbitrary reduction of errors in a single round of the protocol, thereby eliminating the need for concatenation of multiple distillation steps or intermediate fast feedback. Up to a conjecture in the theory of matrix permanents, we show that the resource costs as expressed in the overall number of photons required for the protocol are linear in the required error reduction, and the success probability of the protocol tends to a constant as more photons are used. These two facts render our protocol highly scaleable: to achieve an asymptotic reduction in indistinguishability error by at least a factor of 100 required 59,000 photons with the state-of-the-art scheme, whereas our new scheme only requires 400 photons. We show that our scheme is asymptotically optimal in the degree of error reduction per photon consumed.
These results are based on the study of multiphoton interference in Fourier matrices, using simulation techniques originally designed for boson sampling [9, 21, 22, 23, 24]. We follow up on the observation that the scheme proposed in [17], originally found by numerical optimization, in fact implements a Fourier transform on 3 photons in 3 optical modes. The key difference between interference in small systems (with 3 or 4 photons) and larger system is that in larger systems, multiple herald outcomes are valid for distillation, rendering large distillation schemes much more efficient than expected based on extrapolating the smaller schemes.
Furthermore, we study the application of our protocol in large-scale, fault-tolerant, universal photonic quantum computers, specifically in comparison to quantum error correction (QEC). We find evidence that photon distillation has a higher threshold for distinguishability errors than the surface code. Moreover, we find that a regime exists below the error threshold of the surface code where a combination of photon distillation and QEC is more resource-efficient than only QEC. We discuss how our scheme could operate in practice outside the limit of small errors. These results suggest that an implementation of our photon distillation protocol is perfectly tailored for seamless integration into large-scale photonic quantum computers.
Note added. - Related work was simultaneously released [25].
@@
Figure 1: Artist’s impression of photon distillation. Partially distinguishable photons emitted by imperfect sources are interfered by a photonic distillation circuit. A given set of partial measurements heralds a single photon state with enhanced indistinguishability. After a successful distillation attempt, the distilled photon, with reduced indistinguishability error, is ready for processing in photonic quantum computing.
@@ III Fourier transform-based distillation
In the previous section, we established the framework for imperfect photon interference. To establish intuition for the problem of photon distillation in Fourier matrices, we first discuss the relevant metrics for distillation circuits. We illustrate these metrics by extensively studying the $N=5$ photon Fourier transform with a single photon at each input mode, in both error reduction and heralding probability. A detailed description of our numerical simulation methods is presented in App. B.
The resource cost $C$ of photon distillation schemes is given by the number of photons of a given degree of indistinguishability that is on average required to obtain one photon of a given improved indistinguishability, or equivalently, the degree of improvement that can on average be achieved with a fixed number of photons. This resource cost, in turn, is determined by two factors: the error reduction $R$, i.e. the degree to which a single successful application of the scheme reduces the indistinguishability error on each photon, and the heralding probability $p$, i.e. the probability with which the scheme succeeds in the first place. In the limit of low errors, the error reduction reduces to a multiplicative factor, i.e. $\epsilon^{\prime}=\epsilon/R$, where $\epsilon$ and $\epsilon^{\prime}$ are the degree of error before and after the protocol, respectively. Simultaneously, the associated resource cost for a single distillation step is $C=N/p$.
The current state-of-the-art Fourier transform-based photon distillation protocol is presented in [17], with a degree of error reduction of $R=3$ and a heralding success probability $p=\frac{1}{3}$. Since the available error reduction is limited, further error reduction is available by concatenations, i.e. using the input from one distillation step as the input for the next. However, concatenation has a resource cost quadratic in the overall error reduction required.
Fig. 2 shows the error reduction as a function of input error for the $N=5$ photon Fourier transform-based distillation circuit, for different partial measurement outcomes on modes 1 through 4. Surprisingly, there is more than one valid herald measurement for successful photon distillation. We observe three distinct families of successful heralding patterns.
@@ IV Theoretical analysis
$p=\sum_{j=0}^{N-1}(-1)^{j}(j+1)\prod_{i=1}^{j}\left(1-\frac{i}{N}\right).$ (14)
@@
$p=\frac{\frac{1}{N}\binom{2N-3}{N-1}}{\frac{1}{N}\binom{2N-1}{N}}=\frac{1}{4-\frac{2}{N}}=\frac{1}{4}+\frac{1}{8N}+\mathcal{O}\left(\left(\frac{1}{N}\right)^{2}\right).$ (15)
@@ V Discussion
VI Implementation in photonic quantum computers
We conclude our work by discussing how our findings will find application in large-scale photonic quantum computers. To achieve large-scale quantum computation, errors in the computation must be reduced to a level that on average, much less than one logical error occurs during a computation. Conventionally, this is done by a combination of low intrinsic error rates on the individual components and quantum error correction. The threshold theorem states that for each error correcting code, there exists a threshold $p_{\text{th}}$ such that if $p_{\text{error}}<p_{\text{th}}$, the logical error $p_{\text{L}}$ can be arbitrarily reduced [43]. For the purposes of our discussion, low error rates $p_{\text{error}}$ correspond to better quality photon sources. We address the question of how photon distillation fits in this picture.
First, we discuss the relation between photon distillation and improvements in photon source quality. Photonic quantum computing will require mass manufacturing of photon sources in scalable foundry processes. Therefore, improvement in photon indistinguishability through engineering is ultimately limited by uncontrollable fabrication process errors [44]. Overcoming these limitations by post-fabrication calibration is limited and costly, because accurately identifying the exact cause of partial distinguishability is essential. Both photon distillation and quantum error correction have the advantage that it is not required to identify the exact cause of partial distinguishability. Instead, for photon distillation, the use of quantum interference of single photon wavefunctions provides simultaneous filtering in all internal error modes.
The more direct comparison is therefore with quantum error correction (QEC), where once the engineering of photon sources has reached a natural end point, the choice is to deploy QEC, photon distillation, or a combination of both. There are two regimes of interest: the regime where the probability of error is lower than the error threshold of the error correcting code, and the regime where it is above the threshold. We will work under the assumption that the distinguishability error $\epsilon$ is such that we are operating in the small-error regime of photon distillation. We will see at the end of this section that this assumption is justified for most practically relevant cases, and discuss how our results are modified if it is not.
We will first consider the regime below threshold. Here, the key consideration is resource count: given photon sources of a given quality, and a given target logical error rate, what combination of QEC and photon distillation results in the fewest photon sources required to create a single logical qubit? From the perspective of QEC, partial distinguishability errors are known to cause Pauli stabilizer errors, to reduce the purity of resource states, and to force encoded quantum information out of the code space [45, 16, 46]. In our exploratory analysis of the asymptotically low $\epsilon$ regime, these errors are accounted for in $p_{\text{error}}$. For surface codes, the logical error $p_{\text{L}}$ is reduced by increasing the code distance $d$ [47]:
$p_{\text{L}}=a\cdot\left(\frac{p_{\text{error}}}{p_{\text{th}}}\right)^{\frac{d}{2}},$ (20)
where $a$ is a proportionality constant and $d$ is assumed to be even. The code distance $d$ is a measure for the $d^{2}$ resource states needed to form a surface code patch to encode one logical qubit. Considering a fixed photon number overhead $b$ required to create a resource state, the estimated resource cost for a logical qubit is $C_{\text{L}}=b\cdot d^{2}$ photons.
Linear optical quantum gates can be modeled as the interference of a mixed state of (in)distinguishable photons, followed by measurement and optical switching based on the outcome of that measurement. We make the conservative approximation that these gates fail if even a single one of the injected photons is fully distinguishable. Inspired by Eq. 4, we therefore propose to model $p_{\text{error}}=\mathcal{O}(\epsilon)$  [^5] This is a conservative model: if $p_{\text{error}}=\mathcal{O}(\epsilon^{k})$ for integer $k\geq 2$, the impact of photon distillation will be even greater, as the factor $1/N$ in Eq. 22 will be replaced by $1/N^{k}$.. Photon distillation in the limit of small errors reduces the physical error to $p_{\text{error}}/N$ at the cost of an additional fixed photon number overhead factor of $4N$. Simultaneously, the required code distance to achieve the same logical error as in the scenario without photon distillation is reduced. Specifically, we find that while using photon distillation, the photon resource cost for one logical qubit with similar logical error is
$C_{\text{L}}^{\text{dist.}}=C_{\text{L}}\cdot\left(\frac{2\sqrt{N}\log(\frac{p_{\text{error}}}{p_{\text{th}}})}{\log(\frac{p_{\text{error}}}{p_{\text{th}}})-\log(N)}\right)^{2}.$ (21)
Noticeably, we find that photon distillation outperforms surface code QEC ($C_{\text{L}}^{\text{dist.}}<C_{\text{L}}$) if
$\frac{p_{\text{error}}}{p_{\text{th}}}>\left(\frac{1}{N}\right)^{\frac{1}{2\sqrt{N}–1}}.$ (22)
In Fig. 5, we observe that photon distillation is viable for $p_{\text{error}}/p_{\text{th}}>0.63$, showing that a regime exists where a combination of photon distillation and QEC has reduced resource costs compared to a strategy where only QEC is employed.
Figure 5: Photon distillation resource allocation tradeoff. This figure shows the tradeoff between allocating photon resources for either error mitigation (by means of photon distillation) or enhancing the quantum error-correction code distance. The use of $N$-photon distillation schemes is advantageous when the initial error rate fraction $p_{\text{error}}/p_{\text{th}}$ exceeds the black curve. Generally, the allocation of photon resources to photon distillation is effective when $p_{\text{error}}/p_{\text{th}}>0.63$.
Next, we consider the regime above the QEC threshold. Making a direct comparison between photon distillation and quantum error correction requires establishing a correspondence between logical errors $p_{\textrm{error}}$ and photon errors $\epsilon$. We note that since a photonic quantum computer consumes a fixed number of photons per qubit per computational step, these are both per-qubit, per-step quantities. However, a precise correspondence will depend on the specific architecture used and might even vary depending on where in the computer the error occurs. We note, however, that all photonic quantum gates operate on a small number of $N_{\textrm{gate}}=\mathcal{O}(1)$ photons. Following the approximation that a single error photon induces gate failure, this means that $p_{\textrm{error}}/\epsilon=N_{\textrm{gate}}\approx 1$. This result is consistent with studies of distinguishability errors in specific cases e.g., in dephasing-only error model for cluster states in Ref. [45]. Here, $p_{\text{error}}=\frac{1}{2}-\frac{(1-\epsilon)^{2}}{1+(1-\epsilon)^{2}}=\frac{1}{2}\epsilon+\mathcal{O}(\epsilon^{2})$.
Establishing this correspondence allows us to make a direct comparison: the error threshold for photon distillation scales as $\epsilon^{*}\approx 1/N^{0.61}$, giving results of $\epsilon\simeq 0.2-0.35$, for the regime where Fig. 5 suggests photon distillation is most resource-efficient. In contrast, thresholds for QEC are $p_{\text{th}}=\mathcal{O}(10^{-4}–10^{-1})$ [4, 45], with FBQC having a threshold of up to 1.07% Pauli error [4].
Our results strongly suggest that even for a very pessimistic estimate of $N_{\textrm{gate}}$ of, say, $N_{\textrm{gate}}=10$, there exists a regime where the combination of photon distillation followed by QEC has a higher threshold than only QEC. Furthermore, we note that many existing photon sources operate already below the threshold for photon distillation [25].
Finally, we return to the question of whether the use of the small-error limit is justified in these calculations. For our result below threshold, it is clear that this is the case if we, for example, assume the FBQC threshold of 1.07% error. For the results above the QEC threshold, as noted above, even with a pessimistic assumption about $N_{\textrm{gate}}$, we can show that the theshold of the combined scheme exceeds that of bare QEC. However, a precise calculation of the threshold of a combined photon distillation plus QEC scheme is beyond the present work, as this requires both a better understanding of the high-error regime as well as a better understanding of the relationship between partial distinguishability errors and logical errors in photonic quantum computation. We defer the generalization of including other errors, such as photon loss, for future work.
VII Conclusion
@@
In conclusion, we have used a combination of numerical and theoretical tools to shown that large Fourier transform-based schemes can be used to distill indistinguishable photons from imperfect sources. For this purpose, we explicitly simulated the error reduction for the $N=5$ photon Fourier transform where we showed that there are different paths towards perfect photon indistinguishability. Within the limit of low partial distinguishability error, we have shown modulo a conjecture that the error reduction scales linearly with the number of photons employed in the distillation scheme $N$, while we have given numerical and theoretical evidence that the success probability tends to a constant, giving rise to linear scaling in the number of photons required to achieve a given error reduction.
@@
Nevertheless, applicability of our work strongly depends on the availability of high-fidelity programmable photonic processors and fast optical switches. However, we note that processor size and fidelity of linear optical circuits are constantly improving [30, 49, 50, 51, 52, 53]. Our scalable scheme contributes to the development of adaptive photon distillation circuits, where concatenation strategies are dependent on the actual herald measurements of each distillation step. While current advances in single photon source engineering continue to improve photon indistinguishability [54, 55, 56, 57], we anticipate that photon distillation will aid in the task of creating the millions of almost fully indistinguishable photons necessary for photonic quantum computing. We have argued that a higher error threshold can be achieved by complementing conventional QEC methods with photon distillation, and that a regime exists below threshold where a combination of photon distillation and QEC outperforms bare QEC in terms of resource costs at constant performance. These results are strong evidence that our protocol will find application in large-scale photonic quantum computers.
@@ Appendix C Multi-photon interference coefficients
For comparison, we inspected the $c_{j}$ distribution for Haar-random unitary matrices. We generate $10^{4}$ Haar-random matrices and compute the ratio $\sum_{j=0}^{N}c_{j}j$ over $\sum_{j=0}^{N}c_{j}$ for all mode-assignment lists with exactly one in mode $N$. The histogram of this ratio, for $N=3$ and $N=5$ is plotted in Fig. 10. We observe that the herald measurements for the Fourier transforms (dashed and dotted lines) have extreme ratios $N-1$ that satisfy Eq. 10. These observed extreme ratios are close to the maximum allowed ratio. Since the equality Eq. 9 = Eq. 12 must hold, we require
$\sum_{j=0}^{N}c_{j}-\epsilon\sum_{j=0}^{N}c_{j}j=\sum_{j=0}^{N}c_{j}(1-N\epsilon)+\lambda N\epsilon,$ (28)
where probability $\lambda\geq 0$ to remain physically relevant. Eq. 28 is simplified to find
$\lambda=\sum_{j=0}^{N}c_{j}-\left(\sum_{j=0}^{N}c_{j}j\right)/N.$ (29)
Finally, $\lambda\geq 0$ gives
$N\geq\frac{\sum_{j=0}^{N}c_{j}j}{\sum_{j=0}^{N}c_{j}}.$ (30)
Given that $N\geq N-1$, we conclude that Fourier transforms are optimal in the asymptotically large $N$ regime.